# A Construction of Infinite Families of Self-Orthogonal Quasi-Cyclic Codes Using Constituent Codes

### v2

@@ 2.1. Quasi-Cyclic Codes and Their Representations
$2.$ The Duals of Quasi-Cyclic Codes
@@
    As observed in [HKS, page 136], the irreducible factors $f_{i}$ which are self-reciprocal imply that the cardinality $q_{i}$ of each field extension $F_{i}$ is an even power of $q$, for all ii. However, there are two exceptions. The first one, for all $m$ and $q$, is the field extension coming from the irreducible factor x1x-1 of $x^{m}-1$. The other exception occurs when $q$ is odd and $m$ is even in which $x+1$ is another self-reciprocal factor of $x^{m}-1$. In both exceptions, $q_{i}=q$ and we equip these fields with the Euclidean inner product. Otherwise, we equip $F_{i}$ with the Hermitian inner product.
@@
In the special case when $C_{1}=C_{2}$, the code with generator matrix $G_{1}|G_{1}]$ will be denoted as $2C_{1}$, and such an argument may be naturally extended to $mC_{1}$, for any positive integer $m$. In this work, for any code $C$, we will refer to the code mCmC as $m$-copies of $C$ which will play an important role in the next section.
@@ 3.1.1. Recursive Construction:
for all $1\leq i\leq s-1$. Note that repeated application of (47) yields $d(C_{ui})\geq m^{u-1}d(C_{1i})$ for each ii. This condition can be satisfied, for instance, by setting $C_{ui}=mC_{u-1,i}$. However, unlike in Lemma 3.1, taking copies of constituent codes is not required to obtain our result.
@@
where $\Phi_{d}$ is the $d$-th cyclotomic polynomial, there would be at least four cyclotomic polynomial factors (11, $m$, and $m$’s distinct prime factors). This would result in at least four irreducible factors. If $m$ is prime,
@@ 3.2. Constructing ESO/ESD/EDC QC Codes with Square-Root-Like Lower Bound
$3.$ Galois Closed QC Codes and Their Duals
@@
Let $C\subseteq\mathbb{F}_{q^{2}}^{\ell}$ be a $k$-dimensional code. Then $C$ is Galois closed if and only if $C=\bigoplus_{i}^{k}C_{i}$ where $C_{i}=C_{i}^{q}$ and $\dim(C_{i})=1$, i.e., $C$ is Galois closed if any only if can be decomposed as Galois closed codes of dimension 11.
@@ 4. Constructions of QECCs from QC Codes
Let $q=4$, $\ell=3$, $m=7$, and $\alpha$ be a primitive 77-th root of unity. Let
@@
For any element $\gamma\in\mathbb{F}_{4}(\alpha)\setminus\mathbb{F}_{4}$, let $C_{11}^{\prime}=\left\langle(\gamma,\gamma,\gamma)\right\rangle\subset\mathbb{F}_{4}^{3}(\alpha)$ , $C_{11}^{\prime\prime}=(C_{11}^{\prime})^{\perp}\subset\mathbb{F}_{4}^{3}(\alpha^{3})$, and $C_{11}=\mathbb{F}_{4}^{3}$ be the constituent codes, whose parameters are $3,1,3]_{64}$, $3,2,2]_{64}$, and $3,3,1]_{4}$, respectively. Notice that $C_{11}^{\prime},C_{11}^{\prime\prime}$, and $C_{11}$ are Galois closed by Proposition 3.12, and $C_{11}$ is trivially HDC. Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective 33-QC-code which by Theorem 3.13 is also Galois-closed. The minimum distance of the cyclic code associated to $\mathcal{C}$ and all its associated cyclic subcodes are described below
@@
Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective 33-QC-code. By Theorem 3.7, we have an infinity family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of EDC QC codes whose their parameters are
@@
Let $q=2$, $m=7$, $\ell=8$, and $\alpha$ be a primitive 77-th root of unity. Assume the following decomposition of $R_{2,7}^{8}=\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{7}-1\rangle}\right)^{8}$. Thus
@@
From the minimum distances of the associated cyclic subcodes provided in the Example 4.1, together with the minimum distances of the constituent codes described above, the minimum distance of $\mathcal{C}$ in $\mathbb{F}_{2}^{56}$ is $d\geq d_{GO}=14.$ Moreover, by Lemma 2.1, the dimension of $\mathcal{C}$ is 30, namely, $\mathcal{C}$ is a 88-QC EDC $56,30,\geq 14]_{2}-$code and, based on Corollary 2.11, it is possible to construct a QECC with parameters $56,4,\geq 14]]_{2}$. Notice that in [table], the best $56,4]]_{2}-$QECC has minimum distance 13.

### v3

@@ 2.1. Quasi-Cyclic Codes and Their Representations
2.2. The Duals of Quasi-Cyclic Codes
@@
    As observed in [HKS, page 136], the irreducible factors $f_{i}$ which are self-reciprocal imply that the cardinality $q_{i}$ of each field extension $F_{i}$ is an even power of $q$, for all $i$. However, there are two exceptions. The first one, for all $m$ and $q$, is the field extension coming from the irreducible factor x1x-1 of $x^{m}-1$. The other exception occurs when $q$ is odd and $m$ is even in which $x+1$ is another self-reciprocal factor of $x^{m}-1$. In both exceptions, $q_{i}=q$ and we equip these fields with the Euclidean inner product. Otherwise, we equip $F_{i}$ with the Hermitian inner product.
@@
In the special case when $C_{1}=C_{2}$, the code with generator matrix $G_{1}|G_{1}]$ will be denoted as $2C_{1}$, and such an argument may be naturally extended to $mC_{1}$, for any positive integer $m$. In this work, for any code $C$, we will refer to the code $mC$ as $m$-copies of $C$ which will play an important role in the next section.
@@ 3.1.1. Recursive Construction:
for all $1\leq i\leq s-1$. Note that repeated application of (47) yields $d(C_{ui})\geq m^{u-1}d(C_{1i})$ for each $i$. This condition can be satisfied, for instance, by setting $C_{ui}=mC_{u-1,i}$. However, unlike in Lemma 3.1, taking copies of constituent codes is not required to obtain our result.
@@
where $\Phi_{d}$ is the $d$-th cyclotomic polynomial, there would be at least four cyclotomic polynomial factors ($1$, $m$, and $m$’s distinct prime factors). This would result in at least four irreducible factors. If $m$ is prime,
@@ 3.2. Constructing ESO/ESD/EDC QC Codes with Square-Root-Like Lower Bound
3.3. Galois Closed QC Codes and Their Duals
@@
Let $C\subseteq\mathbb{F}_{q^{2}}^{\ell}$ be a $k$-dimensional code. Then $C$ is Galois closed if and only if $C=\bigoplus_{i}^{k}C_{i}$ where $C_{i}=C_{i}^{q}$ and $\dim(C_{i})=1$, i.e., $C$ is Galois closed if any only if can be decomposed as Galois closed codes of dimension $1$.
@@ 4. Constructions of QECCs from QC Codes
Let $q=4$, $\ell=3$, $m=7$, and $\alpha$ be a primitive $7$-th root of unity. Let
@@
For any element $\gamma\in\mathbb{F}_{4}(\alpha)\setminus\mathbb{F}_{4}$, let $C_{11}^{\prime}=\left\langle(\gamma,\gamma,\gamma)\right\rangle\subset\mathbb{F}_{4}^{3}(\alpha)$ , $C_{11}^{\prime\prime}=(C_{11}^{\prime})^{\perp}\subset\mathbb{F}_{4}^{3}(\alpha^{3})$, and $C_{11}=\mathbb{F}_{4}^{3}$ be the constituent codes, whose parameters are $3,1,3]_{64}$, $3,2,2]_{64}$, and $3,3,1]_{4}$, respectively. Notice that $C_{11}^{\prime},C_{11}^{\prime\prime}$, and $C_{11}$ are Galois closed by Proposition 3.12, and $C_{11}$ is trivially HDC. Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective $3$-QC-code which by Theorem 3.13 is also Galois-closed. The minimum distance of the cyclic code associated to $\mathcal{C}$ and all its associated cyclic subcodes are described below
@@
Let $\mathcal{C}\subset\mathbb{F}_{4}^{m\ell}$ be the respective $3$-QC-code. By Theorem 3.7, we have an infinity family $\{\mathcal{C}_{u}\}_{u\geq 1}$ of EDC QC codes whose their parameters are
@@
Let $q=2$, $m=7$, $\ell=8$, and $\alpha$ be a primitive $7$-th root of unity. Assume the following decomposition of $R_{2,7}^{8}=\left(\frac{\mathbb{F}_{2}[x]}{\langle x^{7}-1\rangle}\right)^{8}$. Thus
@@
From the minimum distances of the associated cyclic subcodes provided in the Example 4.1, together with the minimum distances of the constituent codes described above, the minimum distance of $\mathcal{C}$ in $\mathbb{F}_{2}^{56}$ is $d\geq d_{GO}=14.$ Moreover, by Lemma 2.1, the dimension of $\mathcal{C}$ is 30, namely, $\mathcal{C}$ is a $8$-QC EDC $56,30,\geq 14]_{2}-$code and, based on Corollary 2.11, it is possible to construct a QECC with parameters $56,4,\geq 14]]_{2}$. Notice that in [table], the best $56,4]]_{2}-$QECC has minimum distance 13.